OpenStudy (osuxbucks15):

Round of the measurement to 3 significant figures, 1764.9 mL and 0.007 498 3 times 10 to the 4th power

OpenStudy (osuxbucks15):

@ganeshie8 can you please help I'm very confused

sam (.sam.):

you can think of 1764.9 mL to 3 significant figures as "3 digits" so, you will first simplify to 1.7649x10^3 ml then, since they want 3 digits, \[1.76\times 10^3\] --------------------------------------------------- \[0.0074983=7.4983\times 10^{-3}\] 3 significant figures as "3 digits", \[7.50\times 10^{-3}\]